Let f (x) = x 0 ( u22u+1 )du then find the least value of f (x) in [-1,1].

Let f (x) =
x
0
( u22u+1 )du then find the least value of f (x) in [-1,1].

Grade:12

1 Answers

Sher Mohammad IIT Delhi
10 years ago


its a monotonically increasing function.

f'(x)=(x+1)^2=0, x=-1

minima value=-1/3+1-1=-1/3


Sher Mohammad,
B.tech, IIT Delhi